Find the area of the triangle with vertices at (-3,5),(-3,-5),(5,1)
Answer by HyperBrain(694)

(-3,5) and (-3,-5) are on the same vertical line. Therefore, the base is 10 units
The height is 5+3=8 units
The area is (8)(10)/2=80/2=40 sq. units
Power up,
HyperBrain!
Stop this mess! Click here